Test 2

Determine the maximum power that can be generated from an 18-V emf using any combination of a 6.0-Ohms resistor and a 9.0-Ohms resistor.
1
Expert's answer
2020-04-17T10:52:11-0400

The maximum power that transfers in the circuit

"P_{max}=\frac{V^2}{R_{min}}."

The minimal resistance of circuit would be when two resistors are connected in parallel

"R_{min}=\frac{R_1R_2}{R_1+R_2}=\frac{6\times 9}{6+9}=3.6\:\Omega."

Finally

"P_{max}=\frac{18^2}{3.6}=5\:\rm W."